# Classical JAK2V617F+ Myeloproliferative Neoplasms emergence and development based on real life incidence and mathematical modeling

### v2

@@ Abstract
Mathematical modeling allows us to better understand myeloproliferative neoplasms (MPN) emergence and development. We tested different mathematical models on an initial cohort to determine the emergence and evolution times before diagnosis of JAK2V617F classical MPN (Polycythemia Vera (PV) and Essential Thrombocythemia (ET)). We considered the time before diagnosis as the sum of two independent periods: the time (from embryonic development) for the JAK2V617F mutation to occur, not disappear and enter proliferation, and a second time corresponding to the expansion of the clonal population until diagnosis. We demonstrate using progressively complexified models that the rate of active mutation occurrence cannot be constant, but rather increases exponentially with age following the Gompertz model. We found that the first tumorous cell takes an average time of $63.1\pm 13$ years to appear and start proliferation. On the other hand, the expansion time is constant: $8.8$ years once the mutation has emerged. These results were validated in an external cohort. Using this model, we analyzed JAK2V167F ET versus PV, and obtained that the time of active mutation occurrence for PV takes approximately $1.5$ years more than for ET to develop, while the expansion time was similar. In conclusion, our multi-step approach and the ultimate age-dependent model for the emergence and development of MPN demonstrates that the emergence of a JAKV617F mutation should be linked to an aging mechanism, and indicates a 898-9 years period of time to develop a full MPN.
@@ 1 Introduction
Myeloproliferative Neoplasms (MPN) are blood cancers issued from the transformation of hematopoietic stem cell and related to the acquisition of JAK2 signaling deregulation. The most frequent driver mutation in MPN is JAK2V617F, followed by JAK2 exon12, CALR and MPL [Vainchenker]. MPN are categorized into three main diseases that present the JAK2V617F mutation in the majority of cases: Essential Thrombocythemia (ET), Polycythemia Vera (PV) and Primary Myelofibrosis (PMF) [Tefferi]. ET is characterized by an increased platelet number, PV by erythroid compartment enlargement and PMF by bone marrow fibrosis. When the JAK2 mutation is responsible for the pathology development, the mutation is present in one allele for most cells for ET, but often undergoes an homologous recombination of JAK2V617F for PV resulting in the mutation being present in both alleles [Scott]. In PMF, the JAK2 mutation is often coupled with secondary events such as TET2 mutation or ASXL1 [Vainchenker]. In order to develop one of these pathologies, two steps are needed: the acquisition of a mutation at stem cell level and the development of the cancerous population. When a diagnosis of MPN just like Polycythemia Vera is performed, the total tumor mass (erythrocytes) represents at least $25\%$ of the erythroid hematopoietic mass [Tefferi]. This suggests a total cancer mass of at least 750 grams in the blood and probably the same mass in the bone marrow. Based on the cell mass, we can speculate that a minimum of one million cell divisions are needed to create such a MPN. We also can speculate that the same order or tumor mass is present in all myeloid chronic myeloproliferative malignancies such as ET. Mathematical modelling and statistical analysis through patients data, can help us better understand the dynamics of mutation emergence and expansion, which could yield important information for prevention strategies and early diagnosis.
@@
Two types of data can be found in literature to study the emergence of mutations in HSCs: sequencing data and registries of age of detection. Sequencing data allows to back-track mutations through mathematical models to define the timeline to driver mutation acquisition, (see for example [Osorio2018], [VanEgeren2021], [Mitchell2022]). Detection ages allow to model mutation emergence and expansion as done in [Hermange2022]. In all of these works around MPN emergence, an assumption of a constant mutation occurrence rate throughout the life of an individual was made. Such an Assumption would imply an exponential distribution for the emergence time (see subsection 3.1). However, the available data (Figure 1) shows a growing number of diagnosed patients for older ages (see also [Rohrbacher] and [Moulard]) whose trend is not consistent with this an exponential. In the present work, we considered registries of age of detection. Once the mutation acquisition time was defined, we studied the time until diagnosis, contrary to [Hermange2022] were they studied the behaviour of the size of the mutant population.
Using real life registries of MPN and age of pathology detection in patients, we studied the emergence and expansion of MPN for patients presenting the JAK2V617F mutation. Given the reduced number of reported cases for PMF the data does not allow to properly perform statistical analysis. We then centered our study on ET and PV. We defined the time to diagnosis as composed of two elements: the time $T_{1}$ of active mutation emergence, namely the time for a first cell harboring the JAK2V617F mutation to emerge and enter the cell cycle, and the time $T_{2}$ between emergence and detection, namely the time for cancer population to grow until diagnosis. Then, we presented and tested different models of $T_{1}$ and $T_{2}$ to better define MPN development. The models studied in this work for the age of detection of MPN can be divided into two categories: ones that consider that the active mutation rate is constant as previously seen in literature, and ones that consider it to be dependent on the age of the individual. The first category of models were analysed under different assumptions and we found that they do not seem to be appropriate to explain significantly (from the statistical point of view) the available data. The second type of models, under the assumption that the active mutation occurrence rate grows with age, suggested a much better fit to the data. Hence, we show that in order to properly explain the two available MPN registries, the active mutation occurrence rate rather than being constant, increases exponentially with age, following a Gompertz distribution, conversely to previous models developed.
@@ 2.1 Incidence and frequencies of MPN according to age
In this study, we used data obtained from the cancer registry of the Côte d’Or in France from 1990 to 2020, which compiles the ages of detection for $264$ patients with different types of MPN (Table 1 in Appendix). In this region, life expectancy (and then relative incidence of pathology) can be analyzed. We centered our analysis to patients diagnosed with either PV or TE and presenting the JAK2V617F mutation. The mean detection age in this data set is of 71.9 years with standard deviation of 13.8 years. To further validate the models another set from FIMBANK was used, containing the ages of detection for ET and PV of 1111 individuals as seen in Table 2 in Appendix. In this data set, we have a mean of 64.4 years with a standard deviation of 14.1 years.
The two sets of data used in our analysis were adjusted to take into account the distribution of the population into each age group and consider the proportion of individuals diagnosed within each group. This considers that some individuals could have died before being diagnosed and are thus not present in the data sets. Each frequency was then divided by the probability of survival up to that age group as follows. Let $T_{D}$ be the age of death and $T_{M}$ the age of detection. Then the frequencies given by the data of the probability of being diagnosed at age $t$ is given by
@@
We applied this adjustment based on the French and European survival data by age and sex of 2021 [INED2021]. The resulting corrected data gives us the age distribution of JAK2V617F for patients who have not died of anything else before. This data indicates that the frequencies of JAK2V617F MPN development increases dramatically in older ages as shown in Figure 1. We constructed different models and analyzed their fit with the data.
@@
3 Results
3.1 Modeling mutation occurrence and mutation expansion: a simple model
In order to perform a diagnosis of PV or ET, the total cancer mass needs to surpass a certain threshold of around $M=10^{12}$ cells. We assumed that diagnosis was performed when this size was reached. We consider that the age of detection, $T_{M}$, is the sum of two independent time periods, $T_{1}$ and $T_{2}$: the first one being the time from embryonic development for the mutation to occur, not disappear, and enter in proliferation (that we will name “mutation occurrence”) and a second time corresponding to the expansion time, once the first cancer cell begins to proliferate until the time of MPN diagnosis. The times $T_{1}$ and $T_{2}$ were taken as independent since the time taken for a cell to become proliferative, should not be related to the time for its population to grow to a certain size. In order to consider that the mutation could have appeared at an embryonic level, as suggested in [Mosca], $T_{M}$ was considered to start 9 months before the birth of an individual.
To define the first period of time $T_{1}$, we first assumed that the rate of mutation or activation of an existing mutation at stem cell level is constant through the life of the individual and equal to $\tau$. Then, the time for a mutation to appear follows an exponential distribution with parameter $\tau$. This assumption has been taken in the majority of genetic analysis backtracking reported to date [Osorio2018], [VanEgeren2021], [Mitchell2022]. We considered that each activated mutant cell has a probability $p$ to proliferate and then be detected. The number (1p)(1-p) corresponds to the frequency of non proliferating mutant cells (just like clonal hematopoiesis does for a long period of time) or of mutant cells whose lineages disappear before reaching detection size of MPN. To find the law of $T_{1}$ under these assumptions, let $L$ be a random variable of law $Geom(p)$ that corresponds to the number of mutation events before the first active cancer cell appears and let $X_{1},X_{2},$... be independent and identically distributed random variables of law $Exp(\tau)$ that define the time elapsed between two mutation attempts. Then, $T_{1}$ will simply be the sum of these times:
@@ 3.1 Modeling mutation occurrence and mutation expansion: a simple model
Then, $T_{1}$ is an exponential random variable with parameter $\delta=\tau p$, and its law is given by
$f_{1}(t)=\delta\exp(-\delta t),$
where $t\geq 0$ represents the age of an individual.
@@
The time $T_{2}$ corresponds to the time taken for a population of cells to reach detection size of $10^{12}$ cells. Given the high number of divisions needed to reach this size ($10^{6}$ divisions), variability in division time is averaged and it makes sense to consider that the expansion time should be similar from one individual to the other. The law of $T_{2}$ then needs to be unimodal, centering around a certain value as for example, a lognormal law. The simplest case is to consider a null variance, where all individuals have the same expansion time, implying that environmental modifications do not have an effect on the growth of the tumor. Following this reasoning, we started by assuming that $T_{2}$ was a fixed value $\alpha$ and later studied the case when $T_{2}$ is a lognormal random variable (see Section 3.2).
Then, in our first model the time $T_{M}$ for detection age is distributed as a shifted exponential law, that is for $t\geq\alpha$,
$f_{M}(t)=\delta\exp\left(-\delta\left(t-\alpha\right)\right).$ (1)
The shift $\alpha$ has to be a value between 0 and the minimum age of diagnosis in our cohort. We estimated the parameters $\delta$ and $\alpha$ of the model with a least-squared estimation which minimizes the sum of the square difference between the data and our model for each age. Figure 2 illustrates the data of clinical incidence of MPN in real life compared to the estimation, showing a large disparity between the model and the data. The model follows the opposite trend as the data: it gives a strong probability to small ages, while the data is more focused on older ages. It seems that this model is not appropriate given this type of data. To confirm this, we performed a Chi-square goodness-of-fit test to study the hypothesis that the data follow the distribution of our model and found that it was rejected at a $99\%$ significance level with a p-value smaller than $10^{-17}$, indicating that a constant occurrence rate together with constant expansion time does not fit with the clinical data.
Figure 2: A: Frequency of cases for the data (in black) and probability density function of the estimation (in red) for the classical model (1) B: Accumulated frequencies of cases for the data from the (black) and cumulative distribution of the estimation (in red) for the classical model (1)
3.2 Relaxed assumptions on mutation emergence or expansion time
In an attempt to properly explain the clinical incidence of MPN in the Côte d’Or cohort, we considered two relaxed models that allow additional types of variability: either individual variability on cancer cell emergence or influence of environment on cancer development.
The first model took into account individual variability allowing the time of mutation occurrence to vary from one patient to another or the mutation to develop at different times from one patient to another (quiescence/cycling). This takes into consideration that the biological hypothesis of stress (or environment) influences cancer cell emergence. This model considered again $T_{2}$ as a fixed value $\alpha$ and $T_{1}$ as an exponential random variable, but its rate is different from one patient to another. Then, for $1\leq i\leq N$, where $N$ is the number of patients, we considered that each age of detection $Y_{i}$ is drawn from an exponential distribution with parameter $\delta_{i}$. The $\delta_{i}$’s are considered as an independent sample of a $lognormal(m,s^{2})$, which distribution for $t>0$ is given by
$g_{\delta}(t)=\frac{1}{ts\sqrt{2\pi}}\exp\left(-\frac{\left(\log(t)-m\right)^{2}}{2s^{2}}\right),$
where $t$ represents the age of an individual. This distribution was considered for the $\delta_{i}$’s since it’s of positive support and unimodal. As in the previous model, we obtain that the law of $T_{M}$ is a shifted exponential distribution (1), but the parameter $\delta$ is different from one individual to another. Indeed, for individual $i$, the distribution of the age of detection $T_{M}^{i}$ is given by
$f_{M}^{i}(t)=\exp(-\delta_{i}(t-\alpha)).$ (2)
We can then estimate the values of the parameters $m$, $s$ and $\alpha$ using the software Monolix [Lavielle2014]. This software is well adapted for this mixed effect modeling as it allows to easily estimate the parameters when considering the survival analysis approach. The estimation is based of the SAEM algorithm and gives robust and global convergence. Figure 3 illustrates the data compared to the approximation. The data presents such a big discrepancy that it is sufficient to reject this model.
Figure 3: Accumulated frequencies of cases for the data (in black) and $90\%$ confidence interval of the estimation (in red) for the classical model with individual variability (2).
A second model considered that external factors may not influence the emergence but rather the expansion time and the time of detection. Then, the mutation occurrence rate of JAK2V617F is fixed and constant while the expansion time of cancer development, $T_{2}$ is taken as a random variable. This could be explained by invasion process modifications (ie variation time between two divisions), environment modifications or randomness in the diagnostic process (some patients are diagnosed after vascular events (late detection) when others are diagnosed after a systematic blood cell analysis performed for another pathology (early detection)) or some subjects die (of other unrelated pathologies) before $T_{2}$ was finalized. Considering that it has to be similar between individuals, $T_{2}$ is taken as a lognormal random variable. The distribution of $T_{2}$ is given by
@@ 3.2 Relaxed assumptions on mutation emergence or expansion time
for $t>0$. Since $T_{1}$ and $T_{2}$ are considered independent, we then have that the law of the age of detection $T_{M}$ has the density for any $t>0$
$f_{M}=$ $\int_{0}^{t}f_{1}(s)f_{2}(t-s)ds$ (4)
$=$ $\int_{0}^{t}\delta\exp\left(-\delta s\right)\frac{1}{(t-s)\sigma\sqrt{2\pi}}\exp\left(-\frac{\left(\log(t-s)-\mu\right)}{2\sigma^{2}}\right).$
This model combines an exponential random time for mutation occurrence and a lognormal random time for expansion. The parameters $\delta$, $m$ and $s$ were estimated through a least-squares estimation. This model was also rejected at $99\%$ confidence level through a Chi-square goodness-of-fit test (Figure 4) with a p-value smaller than $10^{-17}$.
Figure 4: A: Frequency of cases for the data (in black) and probability density function of the estimation (in red) for the classical model with lognormal MPN growing time (4). B: Accumulated frequencies of cases for the data (in black) and cumulative density of the estimation (in red) for the classical model with lognormal MPN growing time (4).
The rejection of both of these models imply that even when including individual or environmental variations, models with constant active mutation occurrence $\delta$ cannot explain the data. Thus, alternative models based on an increasing probability of JAK2V617F cancer cell emergence with age were mandatory.
3.3 Age-dependency of JAK2V617F mutation emergence
@@
We consider an active mutation occurrence rate that grows as a person gets older. Biological motivation behind considering an age-dependent mutation occurrence rate can be justified by at least three age-dependent modifications of the environment: i-JAK2V617F mutation appearance (from embryo to adulthood), ii- JAK2V617F activation (ie stem cell quiescence exit) or iii- the success probability of the expansion after activation. Let $\delta(t)$ be the age-dependent active mutation rate, then the distribution of the emergence time $T_{1}$ becomes
$\widehat{f}_{1}(t)=\delta(t)\exp\left(-\int_{0}^{t}\delta(s)ds\right),$ (5)
with $t$ the age of an individual. We study two classical models used in survival analysis for age-dependent rates: the Gompertz model and the Weibull model (see [Lavielle2014], pages 90-92). For an age $t$, the respective functions for the mutation occurrence rate are given by
$\delta_{G}(t)$ $=A\,\exp(k\,t),\quad\text{(Gompertz rate)}$
$\delta_{W}(t)$ $=\frac{\gamma}{\lambda}\left(\frac{t}{\lambda}\right)^{\gamma-1},\quad\text{(Weibull rate)},$
where $A>0$, $k>0$, $\gamma>0$, $\lambda>0$. Since the focus is on studying if age dependency on the active mutation rate can explain the data, we consider the broader model with $T_{2}$ as a lognormal random variable with density function (3). The addition of variability to the expansion time $T_{2}$ should give the same fit or better than a fixed value. We later discuss the case of $T_{2}$ constant. Under these assumptions, the law of $T_{M}$ is given by
$\widehat{f}_{M}(t)=\int_{0}^{t}\widehat{f}_{1}(s)\,f_{2}(t-s)ds,$
since $T_{1}$ and $T_{2}$ are independent, with $\widehat{f}_{1}$ given by (5) and $f_{2}$ given by (3) and $t$ the age of an individual.
For both the Gompertz and Weibull models, we estimated the parameters using the generalized Expectation-Maximization (EM) algorithm (see for example [McLachlan2008]) as presented in [Rai1993].

### v3

@@ Abstract
Mathematical modeling allows us to better understand myeloproliferative neoplasms (MPN), a group of blood cancers, emergence and development. We test different mathematical models on an initial cohort to determine the emergence and evolution times before diagnosis of JAK2V617F+ classical MPN (Polycythemia Vera (PV) and Essential Thrombocythemia (ET)). We consider the time before diagnosis as the sum of two independent periods: the time (from embryonic development) for the JAK2V617F mutation to occur, not disappear and enter proliferation, and a second time corresponding to the expansion of the clonal population until diagnosis. We prove that the rate of active mutation occurrence increases exponentially with age following the Gompertz model rather than being constant. We find that the first tumorous cell takes an average time of $63.1\pm 13$ years to appear and start proliferation. On the other hand, the expansion time is constant: $8.8$ years once the mutation has emerged. These results are validated in an external cohort. Using this model, we analyze JAK2V617F ET versus PV, and obtain that the time of active mutation occurrence for PV takes approximately $1.5$ years more than for ET to develop, while the expansion time was similar. In conclusion, our age-dependent approach for the emergence and development of MPN demonstrates that the emergence of a JAKV617F mutation should be linked to an aging mechanism, and indicates a 898-9 years period of time to develop a full MPN.
@@ 1 Introduction
Myeloproliferative Neoplasms (MPN) are blood cancers issued from the transformation of hematopoietic stem cell and related to the acquisition of JAK2 signaling deregulation. The most frequent driver mutation in MPN is JAK2V617F, followed by JAK2 exon12, CALR and MPL [27]. MPN are categorized into three main diseases that present the JAK2V617F mutation in the majority of cases: Essential Thrombocythemia (ET), Polycythemia Vera (PV) and Primary Myelofibrosis (PMF) [26]. ET is characterized by an increased platelet number, PV by erythroid compartment enlargement and PMF by bone marrow fibrosis. When the JAK2 mutation is responsible for the pathology development, the mutation is present in one allele for most cells for ET, but often undergoes an homologous recombination of JAK2V617F for PV resulting in the mutation being present in both alleles [24]. In PMF, the JAK2 mutation is often coupled with secondary events such as TET2 mutation or ASXL1 [27]. We center our study on ET and PV, given the reduced number of reported cases for PMF. In order to develop one of these pathologies, two steps are needed: the acquisition of a mutation at stem cell level and the development of the cancerous population. The diagnosis of MPN just like Polycythemia Vera is performed when the total tumor mass (erythrocytes) represents at least $25\%$ of the erythroid hematopoietic mass [26]. This suggests a total cancer mass of at least 750 grams in the blood and probably the same mass in the bone marrow. Based on the cell mass, we can speculate that a minimum of one million cell divisions are needed to create such a MPN. We also can speculate that the same order or tumor mass is present in all myeloid chronic myeloproliferative malignancies such as ET.
Mathematical modelling and statistical analysis based on patients data can help us better understand the dynamics of mutation emergence and expansion. This modelling could yield important information for prevention strategies and early diagnosis. Using real life registries of MPN and age of pathology detection in patients, we study the emergence and expansion of MPN for patients presenting the JAK2V617F mutation.
We define the time to diagnosis as composed of two independent elements: the time $T_{1}$ of active mutation emergence, namely the time for a first cell harboring the JAK2V617F mutation to emerge and enter the cell cycle, and the time $T_{2}$ between emergence and detection, namely the time for cancer population to grow until diagnosis. An important parameter in our model is the active mutation rate defined as the product of the mutation rate at stem cell level and the probability of each mutation to develop into a MPN. Based on a first cohort of diagnosis data, we prove here by a rigorous statistical analysis that this active mutation rate is age-dependent and increases exponentially following the Gompertz model (see (4)). We further prove that a constant MPN growing time is sufficient to properly explain the data. The statistical analysis of the model on a second cohort of diagnosis data validates the model and its robustness. The estimation of the parameters of the model allows to obtain information on characteristics of the diseases: an average time of $63.1\pm 13$ years for the first tumorous cell to appear and start proliferation and a constant expansion time of $8.8$ years once the mutation has emerged. We also analyze JAK2V617F ET versus PV, and obtain that the time of active mutation occurrence for PV takes approximately $1.5$ years more than for ET to develop, while the expansion time is similar.
@@
Mathematical modelling of MPN emergence and development based on patients data has already been introduced in different approaches. In [6] the authors use age of detection data to study the behaviour of the size of the mutant population. In other works, sequencing data is used to back-track mutations to define the timeline to driver mutation acquisition, (see for example [18], [28], [13]). In all of these works around MPN emergence, an assumption of a constant mutation occurrence rate throughout the life of an individual is made. Such an assumption would imply an exponential distribution for the emergence time (see subsection 2.2). However, the available data (Figure 1) shows a growing number of diagnosed patients for older ages (see also [22] and [15]) whose trend is not consistent with an exponential (see Figure 4).
In section 2 we describe the data used as well as the construction of the model. Section 3 is dedicated to introducing, selecting and validating our age-dependent model. Finally, in section 4 we apply the selected model to the data to obtain information on the mutation emergence and expansion of MPN and the difference in them for PV and ET.
@@ 2.1 Incidence and frequencies of MPN according to age
In this study, we work with data obtained from the cancer registry of the Côte d’Or in France from 1990 to 2020, which compiles the ages of detection for $264$ patients with different types of MPN (Table 2 in Appendix). In this region, life expectancy (and then relative incidence of pathology) can be analyzed. We center our analysis to patients diagnosed with either PV or TE and presenting the JAK2V617F mutation. The mean detection age in this data set is of 71.9 years with a standard deviation of 13.8 years. To further validate the models, another set from FIMBANK was used, containing the ages of detection for ET and PV of 1111 individuals as seen in Table 3 in Appendix. In this data set, we have a mean of 64.4 years with a standard deviation of 14.1 years.
The two sets of data used in our analysis are adjusted to take into account the distribution of the population into each age group and consider the proportion of individuals diagnosed within each group. This considers that some individuals could have died before being diagnosed and are thus not present in the data sets. Each frequency is then divided by the probability of survival up to that age group as follows. Let $T_{D}$ be the age of death and $T_{M}$ the age of detection. Then the frequencies given by the data of the probability of being diagnosed at age $t$ is given by
@@
We apply this adjustment based on the French and European survival data by age and sex of 2021 [8]. The resulting corrected data gives us the age distribution of JAK2V617F for patients who have not died of anything else before. This data indicates that the frequencies of JAK2V617F MPN development increases dramatically in older ages as shown in Figure 1.
@@
2.2 Model construction and notation
In order to perform a diagnosis of PV or ET, the total cancer mass needs to surpass a certain threshold of around $M=10^{12}$ cells. We assume that diagnosis is performed when this size is reached. We consider that the age of detection, $T_{M}$, is the sum of two independent time periods, $T_{1}$ and $T_{2}$: the first one being the time from embryonic development for the mutation to occur, not disappear, and enter in proliferation (that we will name “mutation occurrence”) and a second time corresponding to the expansion time, once the first cancer cell begins to proliferate until the time of MPN diagnosis. The times $T_{1}$ and $T_{2}$ are taken as independent since the time taken for a cell to become proliferative should not be related to the time for its population to grow to a certain size. In order to consider that the mutation could have appeared at an embryonic level, as suggested in [14], $T_{M}$ is considered to start 9 months before the birth of an individual.
To define the first period of time $T_{1}$, let $\tau$ be rate of mutation or activation of an existing mutation at stem cell level. Then, the time for a mutation to appear follows an exponential distribution with parameter $\tau$. We consider that each activated mutant cell has a probability $p$ to proliferate and then be detected. The number (1p)(1-p) corresponds to the frequency of non proliferating mutant cells (just like clonal hematopoiesis does for a long period of time) or of mutant cells whose lineages disappear before reaching detection size of MPN. To find the law of $T_{1}$ under these assumptions, let $L$ be a random variable of law $Geom(p)$ that corresponds to the number of mutation events before the first active cancer cell appears and let $X_{1},X_{2},$... be independent and identically distributed random variables of law $Exp(\tau)$ that define the time elapsed between two mutation attempts. Then, $T_{1}$ will simply be the sum of these times:
@@ 2.2 Model construction and notation
Then, $T_{1}$ is an exponential random variable with parameter $\delta=\tau p$, which we will refer to as active mutation rate. If we assume that $\delta$ depends on the age $t\geq 0$ of an individual, the density of $T_{1}$ is given by
$f_{1}(t)=\delta(t)\exp\left(-\int_{0}^{t}\delta(s)\,ds\right).$ (1)
@@
On the other hand, the time $T_{2}$ corresponds to the time taken for a population of cells to reach detection size of $10^{12}$ cells. Given the high number of divisions needed to reach this size ($10^{6}$ divisions), variability in division time is averaged and it makes sense to consider that the expansion time should be similar from one individual to the other. The law of $T_{2}$ then needs to be unimodal, centering around a certain value and positive as, for example, a lognormal law whose density for an age $t\geq 0$ is
@@
with $\mu\in\mathbb{R}$, $\sigma>0$. The simplest case is to consider a null variance with $T_{2}=\alpha$ for $\alpha\geq 0$, where all individuals have the same expansion time, implying that environmental modifications do not have an effect on the growth of the tumor.
By independence of $T_{1}$ and $T_{2}$, the law of $T_{M}$ is computed as the convolution of the densities $f_{1}$ and $f_{2}$ of $T_{1}$ and $T_{2}$, respectively. That is, for $t\geq 0$, it is given by
$f_{M}(t)=\int_{0}^{t}f_{1}(s)\,f_{2}(t-s)\,ds.$ (3)
The variables and parameters are summarized in the following table.
Variable Description
$T_{M}$ Time to detection. $T_{M}=T_{1}+T_{2}$
$T_{1}$ Mutation emergence time.
$T_{2}$ Expansion time.
Parameter Description
$\delta$ Active mutation rate. $\delta=\tau p$
$\tau$ Mutation rate.
$p$ Probability of proliferation and detection of a mutant cell.
Table 1: Variables and parameters of the model.
3 Mathematical model and results
3.1 Mathematical age-dependent model
@@
Data in Figure 1 shows a clear age structure, with an increasing number of cases as age increases. Then, we hypothesize an age-dependent active mutation occurrence rate that grows as a person gets older. Biological motivation behind considering an age-dependent mutation occurrence rate can be justified by at least three age-dependent modifications of the environment: i-JAK2V617F mutation appearance (from embryo to adulthood), ii- JAK2V617F activation (ie stem cell quiescence exit) or iii- the success probability of the expansion after activation. We consider the Gompertz model (see [10], pages 90-92) for the active mutation rate $\delta$, which is given by
$\delta_{G}(t)$ $=A\,\exp(k\,t),$ (4)
where $A>0$, $k>0$, and $t$ is the age of an individual. The Gompertz model has been widely used to model different biological age-dependent rates such as mortality [5] or cancer death rate [21]. In our case, it can be explained as follows: we assume that individuals have an initial mutation occurrence rate $A$, and an age-related increase in mutation occurrence represented through a a parameter $k$. Then, for an age $t$, the active mutation occurrence takes the form $\delta_{G}(t)$ and the density of $T_{1}$ in (1) is expressed as
$f_{1}(t)=A\,\exp\left(\frac{A}{k}\right)\,\exp(k\,t)\exp\left(-\frac{A}{k}\,\exp(k\,t)\right)$
We consider two models for the expansion time. The first one being a fixed $T_{2}$ assuming that environmental modifications do not have an effect on the growth of the tumor. The second model takes $T_{2}$ as a lognormal random variable defined by (2). This setting can be explained by invasion process modifications (ie variation time between two divisions), environment modifications or randomness in the diagnostic process (some patients are diagnosed after vascular events (late detection) when others are diagnosed after a systematic blood cell analysis performed for another pathology (early detection)) or some subjects die (of other unrelated pathologies) before $T_{2}$ was finalized. Then, we have two age-dependent models characterized by
Model A.1: $\begin{cases}\delta_{G}(t)=A\,\exp(k\,t)\\ T_{2}=\alpha\text{ fixed},\end{cases}$ (5)
Model A.2: $\begin{cases}\delta_{G}(t)=A\,\exp(k\,t)\\ T_{2}\sim lognormal(\mu,\sigma^{2}).\end{cases}$ (6)
The respective laws, $f_{1}^{M}$ and $f_{2}^{M}$, of the age of detection $T_{M}$ under each model come from (3) and are given for an age $t$ by
$f_{1}^{M}(t)$ $=A\,\exp\left(\frac{A}{k}\right)\,\exp(k(t-\alpha))\exp\left(-\frac{A}{k}\,\exp(k(t-\alpha)\right)$
$f_{2}^{M}(t)$ $=\int_{0}^{t}A\,\exp\left(\frac{A}{k}\right)\,\exp(k\,s)\exp\left(-\frac{A}{k}\,\exp(k\,s)\right)\frac{1}{t\sigma\sqrt{2\pi}}\exp\left(-\frac{\left(\log(t-s)-\mu\right)}{2\sigma^{2}}\right)ds.$
3.2 Parameter estimation